OpenStudy (anonymous):

Solve the system by using the substitution method. x-5y=32 8y-9x=-103

OpenStudy (anonymous):

x-5y=32 x=32+5y Just substitute that equation to 8y-9x=-103

OpenStudy (anonymous):

that makes no sense

OpenStudy (anonymous):

replace x in eq. 8y-9x=-103 with x = 32+5y. Try it first!